an open box is to be constructed from a square piece of sheet metal by removing a square of side 2 feet from each corner and turning up the edges. If the box is to hold 32 cubic feet, what should be the dimensions of the sheet metal?
Answer by mananth(16949)

Let the length of square be x feet
you are cutting off 2 feet from opposite sides.
On a length you are removing 4 feet
The length of the box will be x-4. So the width.
The height will be 2 feet , the length of the cut portion
Volume of cube = L*W*H
(x=4)*(x-4)*2
2*(x-4)^2= 32
(x-4)^2= 16
x-4 = 4
x=8
The square sheet of metal must be 8 feet